OpenStudy (thatonegirl_):

Rule for the nth term of a_16=52, d=5?

rishavraj (rishavraj):

\[a_n = a_1 + (n - 1)d\]

OpenStudy (thatonegirl_):

Can you explain how to find that? I have a bunch like it

rishavraj (rishavraj):

\[a_16 = 52 ~~(given)\] also \[a_{16} = a_1 + (n - 1)d\] so \[52 = a_1 + (16 - 1)d\]

OpenStudy (thatonegirl_):

Is that top formula just the general one????

OpenStudy (thatonegirl_):

and what about d=5?

rishavraj (rishavraj):

the top formula is a general one.... plug d = 5 in the equation

OpenStudy (thatonegirl_):

so \[16=a_1+(n-1)5\]

rishavraj (rishavraj):

nah nah.....it would be \[52 = a_1 + (16 - 1)5\] get a_1 and plug it in general equation for nth term

OpenStudy (thatonegirl_):

Wait so what's the answer then?

OpenStudy (thatonegirl_):

oh nvm I see.. I thought u said that the general equation was the answer lol

rishavraj (rishavraj):

on solving tht equation u will get a_1 = -23 so for nth term \[a_n = a_1 + (n - 1)d\] plug a_1 and d

OpenStudy (thatonegirl_):

\[a_n=-23+(n-1)5\] ?

OpenStudy (thatonegirl_):

Well wouldn't it actually be \[a_n=-28+5n\] ??

rishavraj (rishavraj):

yup....:))

OpenStudy (thatonegirl_):

Cool thank you very much!

rishavraj (rishavraj):

yw